# Inside Out: Externalizing Assumptions in Data Analysis as Validation Checks

### v1

@@ 1 Introduction
In data analysis, analysts often rely on their prior knowledge or domain expertise to form expectations and assess whether results align with them. When results deviate from these expectations, experienced analysts check assumptions made about the data and during the analysis, but often do so without discussing the underlying reasoning in the publications. While publication of analysis code and data is now a common a requirement for the sake of reproducibility [Peng, 2011], the published code alone is often insufficient for understanding the thought process behind an analysis. Code corresponding to published analyses often reflect the final decisions made about analysis and do not reveal the decisions or assumptions made about the data during the analysis. Thus, a reader looking at published code can often be left with many questions about why certain choices were made.
@@
We might gain insight into analysts’ thought processes by speaking with them directly or watching them work via screencast videos they produce, such as, TidyTuesday screencast videos or think-aloud type studies [e.g. Gu et al., 2024]. However, direct observation of analysis is not scalable and may not always be feasible; creating educational screencast videos requires significant effort from the analysts. Ideally, there could be a way to explicitly communciate these throught processes in the data analysis to others. Even better, if these expectations were machine-readable, we could analyze them and learn from the analysis itself. For example, we could answer questions about whether the checks also apply to other researchers analyzing new data in the same context, whether they reflect common practices in the field, or whether they are specific to the data or analysis at hand. Externalizing these thought processes is a practice that has potential to improve the trustworthiness of the analyses [Yu and Barter, 2024] in general and the trustworthiness of subsequent products, such as machine learning models, that may be built on such analyses. Developing an approach to reveal some of this process without requiring a reader to essentially reconstruct the analysis process from the raw data would provide an improved basis for trusting an analysis result [Peng and Hicks, 2021].
@@
In this paper, we conceptualize these inexplicit expectations as analysis validation checks, which allows us to examine the assumptions made about the data during an analysis. We then introduce a procedure to identify the subset of checks that best predict the occurrence of unexpected outcomes. The procedure, based on simulations of the original data, compares the accuracy and independence of different combinations of analysis validation checks. Accuracy is determined using binary classification metrics (precision and recall) from a logic regression model [Ruczinski et al., 2003], while independence is measures the shared information among checks. The proposed workflow offers a numerical guarantee that the analysis will produce the expected results, assuming the assumptions about the data generating mechanism hold.
@@
The rest of the paper is organized as follows: Section 2 reviews the concepts of diagnosing unexpected outcomes and general data quality checks. Section 3 introduces the concept of analysis validation checks, illustrated with a toy example based on step count data. Section 4 describes the procedure that selects the best subset of checks for understanding the occurrence of unexpected outcomes. Section 5 applies this procedure to a larger example that estimates the effect of particulate matter air pollution on daily mortality. Section 6 summarises the paper and discusses a few key considerations.
@@ 2.2 Data analysis checks
The concept of data quality has been studied in the literature, with earlier work focusing on defining frameworks – such as dimensions, attributes, and measures – to improve data quality in database and information systems [Wang and Strong, 1996, Batini et al., 2009, Sidi et al., 2012, Woodall et al., 2014, Cai and Zhu, 2015, Cichy and Rass, 2019]. More recently, data validation has been incorporated into frameworks like Google TensorFlow [Polyzotis et al., 2019] to ensure the quality of data for training machine learning models, as well as for supporting business decision-making [Schelter et al., 2018]. With the increasing availability of open data in scientific research, the users of the data are often not the original data collector and may not be aware of all the details or nuances of the data. This encourages researchers to conduct data quality checks before beginning the analytic process, helping to avoid unexpected discoveries later on. In R, packages like skimr [Waring et al., 2022] and dataMaid [Petersen and Ekstrøm, 2019] provide basic data screening and data quality reports, while packages such as assertr [Fischetti, 2023], validate [van der Loo and de Jonge, 2021], and pointblank [Iannone et al., 2024] focuse on providing infustructures that allow users to define customized data quality checks.
@@ 3.1 A Toy Example
Consider a 30-day step count experiment in a public health setting. Subjects are instructed to walk at least 8,000 steps each day, with an expected average of 9,000 steps, tracked by a step counter app. With data of this nature, we may expect there to be occasional “low” days due to factors such as unfavorable weather conditions limiting outdoor activities. We may also expect “high” days recorded after an outdoor activities or intense workout. Given the requirements of the study, we form our expectation as follows:
@@
To diagnose potential reasons why this expectation might fail, we can establish a few analysis validation checks in anticipation of seeing the data. For example, we can check the quantile of the step count, if more than a third of the days fall below 8,000, or more than a third exceed 10,000 steps, this could indicate an excess of low-count or high-count days. Similarly, we may may expect the standard deviation of the step count not to be overly large. These considerations yield the following three analysis validation checks that fail when:
@@
Figure 1: Beeswarm plot of average step counts across 300 simulated 30-day periods. Each point represents the average step count from one simulation. Similar to a boxplot or violin plot, the beeswarm plot also displays the distribution of each individual data point. The orange points indicate instances where the average step count fails outside the [8,500, 9,500] interval, representing an unexpected outcome in this scenario.
@@ 4 Method
While not all analysis validation checks are equally important, some may be more useful at diagnosing unexpected outcomes than others. This section introduces a procedure to identify a subset of analysis validation checks that best predict the occurence of unexpected outcomes.
The approach uses simulated datasets that have the same structure and characteristics of the observed data to be analyzed in the future. For each simulated dataset, analysis validation checks are applied to record their TRUE or FALSE values. The outcome of the analysis is also computed to determine whether it is unexpected (1) or expected (0). By generating many datasets, each time applying the analysis validation checks and computing the outcome, we can relate the patterns in the checks to the likely that the outcome will be unexpected. Figure 2 provides an overview of the process.
Figure 2: Overview of the procedure to quantify the effectiveness of analysis validation checks. The procedure involves simulating replicates of the data, applying the anlysis, and running the analysis validation checks on each replicate to determine whether the outcome meets the expectation (the unexpected column) and whether each check passes (check 1, check 2, …). The simulation results are then passed to two branches: the accuracy branch calculates the precision and recall of the checks from a logic regression prediction, while the redundancy branch to calculate the mutual information and independence of the checks. These three metrics are combined to produce a single metric that quantifies the effectiveness of the checks in diagnosing unexpected outcomes in the analysis.
@@ 4.1 Accuracy
From the simulated data, the accuracy branch in Figure 2 refers to the set of checks’ ability to detect unexpected outcomes accurately. To relate multiple checks to the outcome, a logic regression model [Ruczinski et al., 2003] is used. Originally developed for SNP microarray data, logic regression constructs Boolean combinations of binary variables, $\mathcal{L}(X_{1},X_{2},\cdots,X_{n})$, to predict classification or regression outcomes. The logical operations allowed to construct $\mathcal{L(\cdot)}$ includes AND ($\land$), OR ($\lor$), and NOT ($\neg$) and an example of the Boolean combination of $\mathcal{L(\cdot)}$ is $X_{1}\text{ AND }(X_{2}\text{ OR }X_{3})$. In our binary-binary classification problem, the logic regression model can be written as follows:
$E(Y)=\mathcal{L}(X_{1},X_{2},\cdots,X_{n})$
where the outcome, $Y$, indicates whether the outcome is unexpected (1) or expected (0), and the analysis validation checks, $X_{1},X_{2},\cdots,X_{n}$, are labeled as 1 if they fail and 0 if they pass.
This objective function is optimized using simulated annealing algorithm to minimize the misclassification error and the following 6 moves are permitted during the optimization to grow or purne the tree (as detailed in Figure 2 of Ruczinski et al. [2003]): 1) replacing a leaf node, 2) replacing an operator, 3) growing a branch, 4) pruning a branch, 5) splitting a leaf node, and 6) deleting a leaf node. Predictions from the logic regression model is then used, along with the observed outcome, to calculate the precision and recall of the optimized Boolean combination of checks. Compared to other tree-based methods for binary-binary prediction, the Boolean combinations from the logic regression model produce a tree structure that can be directly interpreted as the possible combination of checks leading to an unexpected outcome, without the need to invert the tree as required in classic tree-based recursive partitioning methods.
@@ 4.2 Independence
While checks may score high on predictive accuracy, they may be correlated and less useful to provide actions to act on the analysis. This could happen if a set of checks are all tangentially related to the cause of the unexpected results, but none addresses the root cause. The independence score in Figure 2 is defined as 1 - total correlation per observation.
Total correlation is a generalization of mutual information that measures the amount of shared information across multiple variables multiple variables based on Kullback-Leibler distance:
$C(X_{1},X_{2},\cdots,X_{n})=\sum_{x_{1}}\sum_{x_{2}}\cdots\sum_{x_{n}}p(x_{1},x_{2},\cdots,x_{n})\log\frac{p(x_{1},x_{2},\cdots,x_{n})}{p(x_{1})p(x_{2})\cdots p(x_{n})}$
A high total correlation value indicates overlapping of information among the analysis validation checks, while a low value suggests that the checks are independent and provides unique information to diagnose the unexpected outcome. To standardize this measure, the total correlation per observation is used in calculating the independence score.
The three metrics (precision, recall, and independence) can be combiend into a single metric using the arithmetic mean, harmonic mean, or quadratic mean. The differences among these means are minimal when the three metrics are similar. However, as the differences among the metrics increases, the harmonic mean tends to produce the smallest overall score, as it penalizes low values, while the quadratic mean tends to produce the largest score by rewarding higher values more. For simple interpretation of the score, the arithmetic mean is preferred, while in applications where the difference between precision, recall, and independence need to be penalized or rewarded more, the harmonic and quadratic mean should be considered.
@@ 4.3 Toy Example Revisited
Returning to the step count example introduced in Section 3.1, the logic regression model is fitted to the three validation checks described previously to predict the outcome, whether the average number of step falls within the [8,500, 9,500] interval. Figure 3 shows the best-fitting logic regression model as
@@
Logic regression: (check 1 OR check 2) AND (not check 3) 0.861 0.431 0.998 0.669 0.763
Regression tree 0.542 0.780 1.000 0.727 0.774
For comparison, the regression tree produces a similar prediction to the logic regression, however, we argue that the logic regression tree shown in Figure 3 is more interpretable for our purposes because it provides a direct representation of which combinations of analysis checks lead to unexpected outcomes. The logic regression tree is also directly comparable to other diagnostic techniques, which we discuss further in Section 6.
@@ 5 Application
In the study of the health effects of outdoor air pollution, one area of interest is the association between short-term, day-to-day changes in particulate matter air pollution and daily mortality counts. Substantial work has been done to study this question and to date, there appears to be strong evidence of an association between particulate matter less than 10 $\mu$g/m3 in aerodynamic diameter (PM10) and daily mortality from all non-accidental causes [Samet et al., 2000]. For our second example, we use the problem of studying PM10 and mortality along with data from the National Morbidity, Mortality, and Air Pollution Study (NMMAPS) to demonstrate how our analysis validation checks described in Section 4 can be applied. In addition to providing a more substantial problem for our methods, this example also demonstrates how the procedure presented in Section 4 can be used to select cutoff values in the analysis checks to diagnose an unexpected PM10 coefficient from the generalized linear model. The dataset that we use to develop our simulations and analysis checks is from New York City, and contains daily PM10, all-cause (non-accidental) mortality, and average temperature values from 1992–2000. We then will apply the analysis checks to similar data from 39 other cities from the NMMAPS study.
The typical approach to studying the association between PM10 and mortality is to apply a generalized linear model with a Poisson link to relate daily mortality counts to daily measures of PM10. Based on previous work and the range of effect sizes published in the literature, an analyst might expect the coefficient for PM10 in this GLM to lie between $[0,0.005]$, after adjusting for daily temperature [Samet et al., 2000, Welty and Zeger, 2005]. Note that the relatively simple modeling approach being used here is primarily for demonstration; typically far more sophisticated semi-parametric approaches are used in the literature [Peng et al., 2006].
@@
Table 2: A list of checks considered for the generalized linear model of mortality on PM10 and temperature. The checks are based on the sample size, correlation between mortality and PM10, correlation between mortality and temperature, and univariate outlier detection. Multiple cutoff values are specified for each check to determine a failure.
@@
Mortality-PM10 correlation less than $-0.05$
Mortality-PM10 correlation less than $-0.03$
@@
Mortality-temperature correlation greater than $-0.3$
Mortality-temperature correlation greater than $-0.35$
Mortality-temperature correlation greater than $-0.4$
Mortality-temperature correlation greater than $-0.45$
Outlier(s) are presented in the variable PM10
Outlier(s) are presented in the variable mortality

### v2

@@ Abstract
keywords:
diagnostic, logic regression, data analysis assumptions; diagnostic, logic regression, data analysis assumptions
@@ 1 Introduction
In data analysis, analysts often rely on prior knowledge or domain expertise to form expectations and assess whether results align with them. When results deviate from these expectations, experienced analysts check assumptions made about the data during the analysis, but often do so without discussing the underlying reasoning in the publications. While publication of analysis code and data is now a common requirement for the sake of reproducibility [Peng, 2011], the published code alone is often insufficient for understanding the thought process behind an analysis. Code corresponding to published analyses often reflects the final decisions made about analysis and does not reveal the decisions or assumptions about the data during the analysis. Thus, a reader looking at published code can often be left with many questions about why certain choices were made.
@@
We might gain insight into analysts’ thought processes by speaking with them directly or watching them work via screencast videos they produce, such as TidyTuesday screencast videos or think-aloud type studies [e.g. Gu et al., 2024]. However, direct observation of analysis is not scalable and may not always be feasible; creating educational screencast videos requires significant effort from the analysts. Ideally, there could be a way to explicitly communicate these thought processes in the data analysis to others. Even better, if these expectations were machine-readable, we could analyze them and learn from the analysis itself. For example, we could answer questions about whether the checks also apply to other researchers analyzing new data in the same context, whether they reflect common practices in the field, or whether they are specific to the data or analysis at hand. Externalizing these thought processes can improve the trustworthiness of the analyses [Yu and Barter, 2024].
The trustworthiness of an analysis depends not only on the robustness of the methods used but also on a deep understanding of the data. This corresponds to the third step identified in Broderick et al. [2023], where trust can break down if the assumptions underlying an algorithm are not satisfied in practice. In the past, data were often collected directly by analysts through carefully designed experiments, however, contemporary analyses often rely on data curated by government or research institutions, where some data characteristics may be unknown to analysts. When unexpected outcomes arise, they are often later traced back to implicit assumptions made by the analysts - assumptions that were unknowingly violated by the idiosyncrasies in the data. By externalizing these assumptions through analysis validation checks, we aim to make it easier to identify and diagnose such unexpected results.
@@
In this paper, we conceptualize these implicit expectations as analysis validation checks, which allows us to examine the assumptions made about the data during an analysis. We then introduce a procedure to identify the subset of checks that best predict the occurrence of unexpected outcomes. The procedure, based on simulations of the original data, compares the accuracy and independence of different combinations of analysis validation checks. Accuracy is determined using binary classification metrics (precision and recall) from a logic regression model [Ruczinski et al., 2003], while independence measures the shared information among checks. The proposed workflow offers a numerical guarantee that the analysis will produce the expected results, assuming the assumptions about the data generating mechanism hold.
@@
The rest of the paper is organized as follows: Section 2 reviews the concepts of diagnosing unexpected outcomes and general data quality checks. Section 3 introduces the concept of analysis validation checks, illustrated with a toy example based on step count data. Section 4 describes the procedure that selects the best subset of checks for understanding the occurrence of unexpected outcomes. Section 5 applies this procedure to a larger example that estimates the effect of particulate matter air pollution on daily mortality. Section 6 summarizes the paper and discusses a few key considerations.
@@ 2.2 Data analysis checks
The concept of data quality has been studied in the literature, with earlier work focusing on defining frameworks – such as dimensions, attributes, and measures – to improve data quality in database and information systems [Wang and Strong, 1996, Batini et al., 2009, Sidi et al., 2012, Woodall et al., 2014, Cai and Zhu, 2015, Cichy and Rass, 2019]. More recently, data validation has been incorporated into frameworks like Google TensorFlow [Polyzotis et al., 2019] to ensure the quality of data for training machine learning models, as well as for supporting business decision-making [Schelter et al., 2018]. With the increasing availability of open data in scientific research, the users of the data are often not the original data collectors and may not be aware of all the details or nuances of the data. This encourages researchers to conduct data quality checks before beginning the analytic process, helping to avoid unexpected discoveries later on. In R, packages like skimr [Waring et al., 2022] and dataMaid [Petersen and Ekstrøm, 2019] provide basic data screening and data quality reports, while packages such as assertr [Fischetti, 2023], validate [van der Loo and de Jonge, 2021], and pointblank [Iannone et al., 2024] focus on providing infrastructures that allow users to define customized data quality checks.
@@ 3.1 A Toy Example
Consider a 30-day step count experiment in a public health setting. Subjects are instructed to walk at least 8,000 steps each day, with an expected average of 9,000 steps, tracked by a step counter app. With data of this nature, we may expect there to be occasional “low” days due to factors such as unfavorable weather conditions limiting outdoor activities. We may also expect “high” days after outdoor activities or intense workouts. Given the requirements of the study, we form our expectation as follows:
@@
To diagnose potential reasons why this expectation might fail, we can establish a few analysis validation checks in anticipation of seeing the data. [TODO: Add guidelines for designing the checks] For example, we can check the quantile of the step count, if more than a third of the days fall below 8,000, or more than a third exceed 10,000 steps, this could indicate an excess of low-count or high-count days. Similarly, we may expect the standard deviation of the step count not to be overly large. These considerations yield the following three analysis validation checks that fail when:
@@
Figure 1: Beeswarm plot of average step counts across 300 simulated 30-day periods. Each point represents the average step count from one simulation. Similar to a boxplot or violin plot, the beeswarm plot also displays the distribution of each individual data point. The orange points indicate instances where the average step count falls outside the [8,500, 9,500] interval, representing an unexpected outcome in this scenario.
@@ 4 Method
While not all analysis validation checks are equally important, some may be more useful at diagnosing unexpected outcomes than others. This section introduces a procedure to identify a subset of analysis validation checks that best predict the occurrence of unexpected outcomes.
The approach generates simulated datasets from the observed data to compute the outcome and the analysis validation checks (coded as 1 for unexpected and 0 as expected). This then allows us to apply a logic regression tree to link the outcome and the check results to construct a Boolean expression to best predict the unexpected outcomes. Figure 2 provides an overview of the process.
The specific strategy to simulate the data depends on the problem and the nature of the data at hand. Simulated data can be generated using prior domain knowledge, summary statistics derived from the observed data, or statistical techniques such as training-testing splits or bootstrap sampling. While each simulated data may not be independent of the original data - as is the case with bootstrap samples - it should adequately capture the range of possible idiosyncrasies present in the data. Failure to do so can reduce the ability of the resulting Boolean expression to predict unexpected outcomes.
Figure 2: Overview of the procedure to quantify the effectiveness of analysis validation checks. The procedure involves simulating replicates of the data, applying the analysis, and running the analysis validation checks on each replicate to determine whether the outcome meets the expectation (the unexpected column) and whether each check passes (check 1, check 2, …). The simulation results are then passed to two branches: the accuracy branch calculates the precision and recall of the checks from a logic regression prediction, while the redundancy branch calculates the mutual information and independence of the checks. These three metrics are combined to produce a single metric that quantifies the effectiveness of the checks in diagnosing unexpected outcomes in the analysis.
@@ 4.1 Accuracy
From the simulated data, the accuracy branch in Figure 2 refers to the set of checks’ ability to detect unexpected outcomes accurately. To relate multiple checks to the outcome, a logic regression model [Ruczinski et al., 2003] is used. Originally developed for SNP micro-array data, logic regression predicts classification or regression outcomes by constructing Boolean combinations, $\mathcal{L}(X_{1},X_{2},\cdots,X_{k})$, where $X_{i}$ are binary predictors. The logical operations allowed to construct $\mathcal{L(\cdot)}$ include AND ($\land$), OR ($\lor$), and NOT ($\neg$) and an example of the Boolean combination of $\mathcal{L(\cdot)}$ is $X_{1}\text{ AND }(X_{2}\text{ OR }X_{3})$. In our binary-binary classification problem, the logic regression model can be written as follows:
$E(Y)=\mathcal{L}(X_{1},X_{2},\cdots,X_{k}),$
where the outcome, $Y$, indicates whether the outcome is unexpected (1) or expected (0), and the analysis validation checks, $X_{1},X_{2},\cdots,X_{k}$, are labeled as 1 if they fail and 0 if they pass.
This objective function is optimized using a simulated annealing algorithm to minimize the misclassification error and the following six moves are permitted during the optimization to grow or prune the tree (as detailed in Figure 2 of Ruczinski et al. [2003]): 1) replacing a leaf node, 2) replacing an operator, 3) growing a branch, 4) pruning a branch, 5) splitting a leaf node, and 6) deleting a leaf node. Predictions from the logic regression model are then used, along with the observed outcome, to calculate the precision and recall of the optimized Boolean combination of checks. Compared to other tree-based methods for binary-binary prediction, the Boolean combinations from the logic regression model produce a tree structure that can be directly interpreted as the possible combination of checks leading to an unexpected outcome, without the need to invert the tree as required in classic tree-based recursive partitioning methods.
@@ 4.2 Independence
While checks may score high on predictive accuracy, they may be correlated and less useful to provide actions to act on the analysis. This could happen if a set of checks are all tangentially related to the cause of the unexpected results, but none addresses the root cause. To quantify the redundancy among the checks, we use total correlation, $C$, a multivariate generalization of mutual information that captures the amount of shared information across multiple variables. We define the normalized total correlation, $C^{\prime}$, as
$C^{\prime}(X_{1},X_{2},\cdots,X_{k})=\frac{C(X_{1},X_{2},\cdots,X_{k})}{\sum_{i=1}^{k}H(X_{i})}=\frac{\sum_{i=1}^{k}H(X_{i})-H(X_{1},X_{2},\cdots,X_{k})}{\sum_{i=1}^{k}H(X_{i})},$
where $H(X_{i})$ is the entropy of the $i$-th check and $H(X_{1},X_{2},\cdots,X_{k})$ is the joint entropy of checks $\{X_{1},X_{2},\cdots,X_{k}\}$. The amount of independent information, $\eta$, can be defined as:
$\eta(X_{1},X_{2},\cdots,X_{k})=1-C^{\prime}(X_{1},X_{2},\cdots,X_{k})=\frac{H(X_{1},X_{2},\cdots,X_{k})}{\sum_{i=1}^{k}H(X_{i})}.$
To scale this quantity to the range $[0,1]$ for consistency with other metrics (i.e., precision and recall), we define the independence metric as:
$\text{independence}=\frac{\eta-1/k}{1-1/k}=\frac{H(X_{1},X_{2},\cdots,X_{k})/\sum_{i=1}^{k}H(X_{i})-1/k}{1-1/k}.$
An independence metric value of 1 indicates that all checks are independent: $H(X_{1},X_{2},\cdots,X_{k})=\sum_{i=1}^{k}H(X_{i})$ and each check provides unique information to diagnose the unexpected outcome (the trivial case is when there is only one check). A value of 0 indicates the checks are all identical: $H(X_{1},X_{2},\cdots,X_{k})-\sum_{i=1}^{k}H(X_{i})=H(X_{i})/kH(X_{i})=1/k$.
The three metrics (precision, recall, and independence) can be combined into a single metric using the arithmetic mean, harmonic mean, or quadratic mean. The differences among these means are minimal when the three metrics are similar. However, as the differences among the metrics increases, the harmonic mean tends to produce the smallest overall score, as it penalizes low values, while the quadratic mean tends to produce the largest score by rewarding higher values more. For simple interpretation of the score, the arithmetic mean is preferred, while in applications where the difference between precision, recall, and independence need to be penalized or rewarded more, the harmonic and quadratic mean should be considered.
@@ 4.3 Toy Example Revisited
Returning to the step count example introduced in Section 3.1, the logic regression model is fitted to the three validation checks described previously to predict the outcome, whether the average number of steps fall within the [8,500, 9,500] interval. Figure 3 shows the best-fitting logic regression model as
@@
Logic regression: (check 1 OR check 2) AND (not check 3) 0.861 0.431 0.818 0.637 0.703
Comparison: (check 1) AND (not check 3) 0.278 0.870 0.881 0.510 0.676
Regression tree 0.542 0.780 0.818 0.689 0.713
For comparison, we include a simplified rule, (quantile(step, 0.6) > 10,000) AND (NOT sd(step) > 2,500). Although the checks on the 60th quantile and standard deviation are less correlated – reflected in a higher independence score – this pair is less accurate at predicting the unexpected outcome, resulting in lower harmonic and arithmetic mean scores. The regression tree produces a similar prediction to the logic regression, however, we argue that the logic regression tree shown in Figure 3 is more interpretable for our purposes because it provides a direct representation of which combinations of analysis checks lead to unexpected outcomes. The logic regression tree is also directly comparable to other diagnostic techniques, which we discuss further in Section 6.
@@ 5 Application
[…]
The typical approach to studying the association between PM10 and mortality is to apply a generalized linear model (Poisson family with a log link) to relate daily mortality counts to daily measures of PM10. Based on previous work and the range of effect sizes published in the literature, an analyst might expect the coefficient for PM10 in this GLM to lie between $[0,0.005]$, after adjusting for daily temperature [Samet et al., 2000, Welty and Zeger, 2005]. Note that the relatively simple modeling approach being used here is primarily for demonstration; typically far more sophisticated semi-parametric approaches are used in the literature [Peng et al., 2006].
@@
Table 2: List of checks considered for the generalized linear model of mortality on PM10 and temperature. The checks are based on the sample size, correlation between mortality and PM10, correlation between mortality and temperature, and univariate outlier detection. Multiple cutoff values are specified for each check to determine a failure.
@@
Mortality-PM10 correlation less than 0.05-0.05
Mortality-PM10 correlation less than 0.03-0.03
@@
Mortality-temperature correlation greater than 0.3-0.3
Mortality-temperature correlation greater than 0.35-0.35
Mortality-temperature correlation greater than 0.4-0.4
Mortality-temperature correlation greater than 0.45-0.45
Outliers are present in the variable PM10
Outliers are present in the variable mortality